Grade Upto college levelGeneral Physics

suppose potential energy between electron n proton at distance r is given by... (-ke^2)/(3.r^3)...use bohr's theory to obtain energy levels of such hypothetical atom

XYZ Person
15 Years agoGrade Upto college level

1 Answer

Chetan Mandayam Nayakar
15 Years ago

Force=-dV/dr =(ke^2)/r^(-4)=mv^2/r,mvr=nh/2pi, solving the system of equations will yield the answer